Find the equation of the line that is parallel to the graph of y=3x-2 and has the y-intercept (0,3)?
Answer by sehar(13)

the given line is y= 3x-2
let the required line be y=mx+c, where m is slope
slope of given line is (-coeficient of x/coefficient of y) i.e. -3
parallel lines hv equal slopes always so put m=-3 in y=mx+c
y=-3x+c.....(i)
now as (0,3) pass through the reuired line ,it must satisfy (i)
3=0+c
so, c=3
put it in (i)
y=-3x+3
which is the required equation